# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v2

@@ E.1 Toy dataset on 2D and 3D
E.2 Retrieval experiments on CIFAR10 and CUB-200
@@
E.3 Classification on CIFAR-100
For teacher models, we use pretrained models released in the official code of the work Tian et al. (2020) without any modification. We also use the same student models from the same code. We follow the same training scheme as the work Tian et al. (2020). More precisely, for optimizer, we use SGD(lr=0.01, momentum=0.9, weight_decay=5e-4). We use the training set for transferring, during 240 epochs, where we decay the learning by a factor of 0.1 at epochs 150, 180 and 210. The batch size is 64. Data augmentation techniques include random cropping (32x32, padding=4) and random horizontal flipping. We fix temperature $\tau=0.2$ for teacher model and $\tau=0.3$ for student models. Our technique is applied on the penultimate layer and the logit layer (before softmax function). The total loss for training the classification student model is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-entropy}}+\lambda\mathcal{L}_{\text{ours}}$, where we fix $\lambda=5$. We perform 5 independent runs to report results in Table 3. The running time for one run (240 epochs) is as follows: ShuffleNetV1: about 45 minutes. ShuffleNetV2: about 52 minutes. MobileNetV2: about 1 hour.
E.4 Ablation study: effect of scaling temperature in the soft ranking
@@
  • Summary. In practice, intermediate values $\tau\in\{0.1,0.3,0.5\}$ yield the best results. This confirms the need for careful temperature tuning.

### v3

@@ 5.6 Ablation study: Student model size
6 Scope and Practical Considerations
Scope. Our method is intentionally designed for heterogeneous teacher–student settings, where the teacher and student may differ substantially in architecture and capacity. This scenario reflects practical deployment needs, in which knowledge is often transferred from a large, high-performance model to a smaller and structurally different model for efficiency or hardware constraints. In contrast, we are convinced that homogeneous settings (e.g., similar architectures) are typically well served by standard knowledge distillation (KD) approaches. Our goal is therefore not to replace or improve conventional KD in homogeneous cases, but rather to focus on practically relevant heterogeneous regime.
Moreover, our focus is to introduce a new measure, termed perception coherence, for feature representation transfer between models, ideally to enhance representation alignment specifically within the KD setting. Accordingly, our empirical comparisons are restricted to distillation-based baselines. Broader general-purpose representation learning frameworks, such as self-supervised representation learning, are beyond the scope of this study. Nevertheless, these methods are conceptually complementary to our framework, and their integration with the proposed approach represents a promising avenue for future investigation.
Computational Complexity. The theoretical per-batch complexity of the proposed loss is $\mathcal{O}(B^{3})$ due to the computation and comparison of pairwise dissimilarities under soft ranking. However, the implementation relies entirely on batched tensor operations (distance matrix computation, sigmoid-based soft ranking, and squared loss), which are highly optimized on modern GPUs. Moreover, soft-ranking computations for different reference (anchor) points are independent and can be parallelized. With $N$ GPUs, the batch can be partitioned along the reference dimension, allowing each device to process approximately $B/N$ anchors. The final loss is aggregated via standard reduction operations (e.g., all-reduce), enabling near-linear scaling in practice and substantially reducing wall-clock training time.
Besides, it is worth emphasizing that, in many practical and industrial scenarios, lightweight models are trained on servers or high-performance hardware and subsequently deployed to edge devices for inference. Therefore, the training cost does not impact the deployment of the final student model on the edge devices.
Practical Guidance for the Temperature Parameter $\tau$. The temperature parameter $\tau$ in the soft-ranking formulation controls the smoothness of relative distance comparisons. Empirically, we find that values in the range $\{0.1,0.2,0.3\}$ work well for both teacher and student models. From a training-dynamics perspective, using a slightly larger $\tau$ for the student can facilitate optimization, as excessively small $\tau$ values may weaken gradient signals. Conversely, a smaller $\tau$ for the teacher encourages sharper relative ranking, since large $\tau$ values render the sigmoid function more linear and reduce ranking meaning [^5] See Appendix E.5 for a brief analysis about the temperature.. In our experiments, we intentionally use the same $\tau$ across different teacher–student pairs for simplicity and fairness. Nevertheless, task-specific tuning of $\tau$ may further improve performance in practical applications.
@@ E.1 Toy dataset on 2D and 3D
E.2 Stylized Empirical Study of Perception Coherence and Downstream Classification
For simplicity, the teacher and student models share the same architecture, consisting of a feature extractor followed by a classification head (softmax layer). The feature extractor has the following structure $\mathrm{Linear}(2,20)\rightarrow\mathrm{ReLU}\rightarrow\mathrm{Linear}(20,20)$. The classification head (softmax layer) consists of a linear layer $\mathrm{Linear}(20,2)$, which maps the extracted features to two output classes, followed by a softmax function.
The teacher model is first trained with the training set (with binary labels) for 200 epochs using the Adam optimizer with a learning rate of $10^{-3}$. Knowledge transfer is then performed by optimizing the perception coherence of the student model to match the feature structure produced by the teacher’s feature extractor. This is done by using the same training set (no label). Note that during this transfer stage, only the student feature extractor is optimized, while the classification head is excluded. The student model is trained using the Adam optimizer with a learning rate of $10^{-3}$ for 40 epochs. We save the student model every 4 epochs, resulting in a sequence of checkpoints $\{S_{k}\}_{k=1}^{10}$.
For each checkpoint $\{S_{k}\}_{k=1}^{10}$, we freeze the feature extractor and train a new classification head in a supervised manner using the training set (with binary labels). This supervised training is conducted for 20 epochs with the Adam optimizer and a learning rate of $10^{-3}$. Finally, using the test dataset, the test accuracy after this supervised training is recorded for each checkpoint, yielding $\{\mathrm{Acc}_{k}\}_{k=1}^{10}$.
E.3 Retrieval experiments on CIFAR10 and CUB-200
@@
E.4 Classification on CIFAR-100
For teacher models, we use pretrained models released in the official code of the work Tian et al. (2020) without any modification. We also use the same student models from the same code. We follow the same training scheme as the work Tian et al. (2020). More precisely, for optimizer, we use SGD(lr=0.01, momentum=0.9, weight_decay=5e-4). We use the training set for transferring, during 240 epochs, where we decay the learning by a factor of 0.1 at epochs 150, 180 and 210. The batch size is 64. Data augmentation techniques include random cropping (32x32, padding=4) and random horizontal flipping. The most recent VRM method (Zhang et al., 2025) introduces virtual relation matching and applies RandAug (Random Augmentation). For a fair comparison, in addition to random cropping and random horizontal flipping, we also add RandAug to the inputs using the official code associated with this last work. We fix temperature $\tau=0.2$ for teacher model and $\tau=0.3$ for student models. Our technique is applied on the penultimate layer and the logit layer (before softmax function). The total loss for training the classification student model is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-entropy}}+\lambda\mathcal{L}_{\text{ours}}$, where we fix $\lambda=5$. We perform 5 independent runs to report results in Table 4. The running time for one run (240 epochs) is as follows: ShuffleNetV1: about 45 minutes. ShuffleNetV2: about 52 minutes. MobileNetV2: about 1 hour.
E.5 Ablation study: effect of scaling temperature in the soft ranking
@@
  • Summary. In practice, intermediate values $\tau\in\{0.1,0.3,0.5\}$ yield the best results. This suggests that further temperature tuning might be useful to have optimal results.
Appendix F Computational Complexity Comparison
In this section, we compare the computational complexity of different teacher-student pairs in the experiments. As required in practical deployment scenarios, all student models are strictly smaller and computationally cheaper than their corresponding teacher models, with details reports in Table 8. These results are obtained using the python package ptflops (Sovrasov, 2018).
Table 8: Computational complexity comparison (in MACs) of teacher-student pairs. MAC denotes multiply-accumulate operations, where 11 MAC $\approx 2$ FLOPs (one multiplication and one addition).
CIFAR10 SVHN CIFAR100 (ResNet-50 $\rightarrow$ MobileNetV2) CIFAR100 (ResNet-32x4 $\rightarrow$ ShuffleNetV1) CIFAR100 (ResNet-32x4 $\rightarrow$ ShuffleNetV2)
Teacher model 557.22 MMAC 1.82 GMAC 1.31 GMAC 1.09 GMAC 1.09 GMAC
Student model 523.15 KMAC 37.26 MMAC 7.35 MMAC 41.58 MMAC 46.06 MMAC